Problem 1-6 The rate of return on Cherry Jalopies, Inc., stock over the last five years was 16 percent, 11 percent, -1 percent, 6 percent, and 11 percent. Over the same period, the return on Straw Construction Companys stock was 16 percent, 23 percent, -6 percent, 6 percent, and 11 percent. Calculate the variances and the standard deviations for Cherry and Straw. (Enter variance as a decimal and standard deviation as a percent. Round your variance to 5 decimal places and standard deviation to 2 decimal places. Omit the% sign in your response.) Cherry Straw Variance Standard deviation
Answer #1

Cherry:

Average return=Total return/Total time period

=(16+11-1+6+11)/5=8.6%

Return (Return-Average return)^2
16 (16-8.6)^2=54.76
11 (11-8.6)^2=5.76
-1 (-1-8.6)^2=92.16
6 (6-8.6)^2=6.76
11 (11-8.6)^2=5.76
Total=165.2%


Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[165.2/(5-1)]^(1/2)

=[165.2/4]^(1/2)

=6.43%(Approx)

Variance=Standard deviation^2

=0.00413

Straw:

Average return=Total return/Total time period

=(16+23-6+6+11)/5=10%

Return (Return-Average return)^2
16 (16-10)^2=36
23 (23-10)^2=169
-6 (-6-10)^2=256
6 (6-10)^2=16
11 (11-810)^2=1
Total=478%


Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[478/(5-1)]^(1/2)

=[478/4]^(1/2)

=10.93%(Approx)

Variance=Standard deviation^2

=0.01195

Cherry Straw
Variance 0.00413 0.01195
Standard deviation 6.43 10.93(Approx).